# Context-aware Decoding Reduces Hallucination in Query-focused Summarization

### 21 Dec 2023 v1

@@ Abstract
Query-focused summarization (QFS) aims to provide a summary of a single document/multi documents that can satisfy the information needs of a given query. It is useful for various real-world applications, such as abstractive snippet generation or more recent retrieval augmented generation (RAG). A prototypical QFS pipeline consists of a retriever (sparse or dense retrieval) and a generator (usually a large language model). However, applying large language models (LLM) potentially leads to hallucinations, especially when the evidence contradicts the prior belief of LLMs. There has been growing interest in developing new decoding methods to improve generation quality and reduce hallucination. In this work, we conduct a large-scale reproducibility on one recently proposed decoding method—Context-aware Decoding (CAD). In addition to replicating CAD’s experiments on news summarization datasets, we include experiments on QFS datasets, and conduct more rigorous analysis on computational complexity and hyperparameter sensitivity. Experiments with eight different language models show that performance-wise, CAD improves QFS quality by (1) reducing factuality errors/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores, while also at a cost of increased inference-time FLOPs and reduced decoding speed. The code implementation based on Huggingface Library is made available  [^1] https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs..
@@ 1 Introduction
In this work, we focus on Context-aware Decoding [Shi+23] (CAD). CAD leverages the idea of pointwise mutual information and proposes a product-of-experts enhancement to the vanilla decoding method to make the generation more conditioned on the input evidence. We conduct a large-scale reproducibility with different choices of language models to study CAD’s effectiveness in query-focused summarization, and include more rigorous analysis on computational complexity and hyperparameter sensitivity. Extensive experiments show that CAD improves QFS quality by (1) reducing factual mistakes/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores. On the other hand, CAD also introduces additional inference-time FLOPs and potentially slows down decoding speed.
@@ 5 Related Work
[Ji+23, ] present a survey of hallucination in natural language generation. There are mainly two types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. Intrinsic hallucination refers to the generated output that contradicts the source content, whereas extrinsic hallucination refers to the generated output that cannot be verified from the source content. In retrieval augmented generation, we mainly focus on intrinsic hallucination. Pre-trained language models (not instruction finetuned) are more prone to hallucination, as the gigantic pre-training corpus may contain untruthful texts and the teacher forcing training procedure lets the models pick up this untruthful information. [Shu+21, ] show that retrieval augmentation can reduce hallucinations in downstream tasks. [Tia+23, ] show that language models can be trained to reduce hallucination via reinforcement learning techniques.

### 31 Dec 2023 v2

@@ Abstract
Query-focused summarization (QFS) aims to provide a summary of a single document/multi documents that can satisfy the information needs of a given query. It is useful for various real-world applications, such as abstractive snippet generation or more recent retrieval augmented generation (RAG). A prototypical QFS pipeline consists of a retriever (sparse or dense retrieval) and a generator (usually a large language model). However, applying large language models (LLM) potentially leads to hallucinations, especially when the evidence contradicts the prior belief of LLMs. There has been growing interest in developing new decoding methods to improve generation quality and reduce hallucination. In this work, we conduct a large-scale reproducibility study on one recently proposed decoding method—Context-aware Decoding (CAD). In addition to replicating CAD’s experiments on news summarization datasets, we include experiments on QFS datasets, and conduct more rigorous analysis on computational complexity and hyperparameter sensitivity. Experiments with eight different language models show that performance-wise, CAD improves QFS quality by (1) reducing factuality errors/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores, while also at a cost of increased inference-time FLOPs and reduced decoding speed. The code implementation based on Huggingface Library is made available  [^1] https://github.com/zhichaoxu-shufe/context-aware-decoding-qfs..
@@ 1 Introduction
In this work, we focus on Context-aware Decoding [Shi+23] (CAD). CAD leverages the idea of pointwise mutual information and proposes a product-of-experts enhancement to the vanilla decoding method to make the generation more conditioned on the input evidence. We conduct a large-scale reproducibility study with different choices of language models to study CAD’s effectiveness in query-focused summarization, and include more rigorous analysis on computational complexity and hyperparameter sensitivity. Extensive experiments show that CAD improves QFS quality by (1) reducing factual mistakes/hallucinations while (2) mostly retaining the match of lexical patterns, measured by ROUGE scores. On the other hand, CAD also introduces additional inference-time FLOPs and potentially slows down decoding speed.
@@ 5 Related Work
Search as an Information-seeking Process.
[Mar06] categorize search into three two different information seeking processes—Information Lookup and Exploratory Search. Information lookup refers to the user looking for simple facts or asking factoid questions, while exploratory search/search to learn often involves multiple iterations. Different approaches have been explored to reduce users’ cognitive loads in the information-seeking process. For example, Perplexity AI [^3] https://www.perplexity.ai/ transfers web search queries into QA-style questions, and cites sources in the answers to improve verifiability. Snippets [BC14] provide short previews for each webpage to allow users to locate relevant information without examining the webpage. Counterfactual explanations [Xu+23] suggest counterfactual queries as explanations and demonstrate search sessions’ effectiveness can be improved.
@@
[Ji+23, ] present a survey of hallucination in natural language generation. There are mainly two types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. Intrinsic hallucination refers to the generated output that contradicts the source content, whereas extrinsic hallucination refers to the generated output that cannot be verified from the source content. In retrieval augmented generation, we mainly focus on intrinsic hallucination. Pre-trained language models (not instruction finetuned) are more prone to hallucination, as the gigantic pre-training corpus may contain untruthful texts and the teacher forcing training procedure lets the models pick up this untruthful information. [Shu+21, ] show that retrieval augmentation can reduce hallucinations in downstream tasks. [Tia+23, ] show that language models can be trained to reduce hallucination via reinforcement learning techniques. [Var+23, ] propose to use lm logits to detect hallucination in the generation process and mitigate it by retrieving evidence and rewriting.